Help!Linear systems wasn't bad enough now it's linear word problems. I'm totally lost. I would appreciate any help you can be. Thank-you ahead of time you generous soul. Orelan Paint Town sold 45 paintbrushes, one
kind at $8.50 each and another at $9.75 each. In all,
$398.75 was taken in for the brushes. How many of
each kind were sold?


Answer by stanbon(75887)

Orelan Paint Town sold 45 paintbrushes, one
kind at $8.50 each and another at $9.75 each. In all,
$398.75 was taken in for the brushes. How many of
each kind were sold?
------------------
Quantity Equation: a + b = 45
Value Equations: 850a+ 975b = 39875 cents
=============================================
Multiply thru the 1st by 850 to set it up for elimination:
850a + 850b = 850*45
850a + 975b = 39875
------------------------------
Subtract the 1st Eq. from the 2nd so you can solve for "b":
125b = 1625
b = 13 (there are 13 of the $9.75 pens)
--------------
Substitute into a + b = 45 to solve for "a":
a + 13 = 45
a = 32 (there are 32 of the $8.50 pens)
===============================================
Cheers,
Stan H.